OpenStudy (anonymous):

Kobe's overtime pay is $5 an hour more than his regular pay. He worked 8 hours at his regular wage and 3 hours at his overtime wage. He earned $114. What is Kobe's regular wage per hour?

OpenStudy (jojokiw3):

So you have 2 unknowns, this means you need 2 equations to solve 2 unknowns, what 2 equations do you have?

OpenStudy (anonymous):

No equations are given.

OpenStudy (jojokiw3):

Well 2 algebraic equations are in the word problem, you just gotta convert it.

OpenStudy (anonymous):

Oh. 8x + 3(5) = 114 is the only equation I can come up with.

OpenStudy (jojokiw3):

Hm, Let's see again. Let's name Overtime Pay as "O" and Regular Pay as "R". What does O equal in terms of R?

OpenStudy (anonymous):

?

OpenStudy (jojokiw3):

? What does Overtime equal? I don't want a number, because you don't know it and neither do I. The word problem tells you what his overtime pay equals.

OpenStudy (anonymous):

More time than the assigned hours

OpenStudy (jojokiw3):

My bad. what does his overtime PAY equal?

OpenStudy (anonymous):

The amount of money you're given when you work more hours than you're supposed to.

OpenStudy (jojokiw3):

Equal, not definition.

OpenStudy (anonymous):

?

OpenStudy (jojokiw3):

It says in the first sentence of the problem

OpenStudy (jojokiw3):

It tells you how much his overtime pay is.

OpenStudy (anonymous):

Kobe's overtime pay is $5 an hour more than his regular pay. Oh.

OpenStudy (jojokiw3):

Yeah. Make that an equation.

OpenStudy (anonymous):

x = x + 5

OpenStudy (jojokiw3):

How can x = x+5?

OpenStudy (jojokiw3):

Okay so you don't know what his overtime pay is, or what his regular pay is. These are called unknowns and you have 2 of them. You have to assign unique letters for each unknown.

OpenStudy (anonymous):

Sorry. Um... I don't know. Sorry I don't like math so I don't understand a lot of things.

OpenStudy (jojokiw3):

It's okay, we all have things we're learning.

OpenStudy (jojokiw3):

Okay so when you say "x = x+5" you're saying some thing like, the amount of my apples is the same as the amount of my apples PLUSES five more apples. That doesn't work right?

OpenStudy (anonymous):

I guess you're right.

OpenStudy (jojokiw3):

So once you used x for "overtime pay", you can't use it again for regular pay. Instead, use another letter.

OpenStudy (anonymous):

Oh how about y.

OpenStudy (jojokiw3):

Yeah, that works! But now that you've decided on x and y, they'll mean the same thing for the rest of the problem. So now you have x = y+5, right?

OpenStudy (anonymous):

Yeah!

OpenStudy (jojokiw3):

So we've gotten everything out of the first sentence, next two sentences. "He worked 8 hours at his regular wage and 3 hours at his overtime wage. He earned $114."

OpenStudy (jojokiw3):

See if you can make an equation out of that. Remember, x = overtime pay and y = regular pay.

OpenStudy (anonymous):

Okay so this is the part where I get really confused. I tried 3(5) so that it represents : the amount of money he got(5) for working extra hours(3) and I got 15 and I did 114 - 15 but I got 99 and 99 is not one of the choices and I don't know what I did wrong.

OpenStudy (jojokiw3):

Haha okay, holld them reins. You still don't know what x or y is, and you need to find those out first.

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

Sorry.

OpenStudy (jojokiw3):

y can be 1000, 10000, 1 mil, we have to figure what y is in order to figure what x is.

OpenStudy (jojokiw3):

Now lemme show you the equation for those two sentences.

OpenStudy (anonymous):

Okay.

OpenStudy (jojokiw3):

\[8(y) + 3(x) = 144\]

OpenStudy (jojokiw3):

Do you have any questions?

OpenStudy (anonymous):

Um...You wrote 144 but its 114.

OpenStudy (jojokiw3):

Hahah! Whooppps. See, I'm human too. \[8y + 3y = 114\]

OpenStudy (anonymous):

Haha yeah well that's okay. Well thanks anyway! Wow you're great at math, how do you figure that all out?

OpenStudy (jojokiw3):

OHoafhohodf. UGH another mistake. No nononono. Get rid of that equation oh god.

OpenStudy (jojokiw3):

\[8y + 3x = 114\]

OpenStudy (jojokiw3):

There, that's correct now. whew.

OpenStudy (anonymous):

Okay. Thanks! :)

OpenStudy (jojokiw3):

Can you solve from there?

OpenStudy (anonymous):

Yeah! Thank You, Thank You and Thank You a million more times. Bye!

OpenStudy (jojokiw3):

Haha alright. Good luck. :)

OpenStudy (anonymous):

Thanks!